# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v2

@@ Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into functional code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, specifically from the original paper authors, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
@@ 1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge [7, 3, 32]. However, across many disciplines, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only 21.23 % of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of scientific innovation.
@@
To answer this question, we introduce PaperCoder, a multi-agent LLM-powered framework, designed to automatically generate executable code repositories in machine learning directly from and contextualized with research papers, which differs from prior work that requires partial implementations from human inputs. Specifically, PaperCoder aims to emulate the typical life cycle of human developers and researchers in writing the repository-level code, by decomposing the task into three structured stages: planning, analysis, and generation. First, during the planning stage, the proposed framework constructs a high-level roadmap to identify core components to implement, draws the overall system architecture with class and sequence diagrams to model structural relationships between modules, identifies file dependencies with their execution orders to guide correct build and execution flows, and generates configuration files to enable flexible customization of experimental workflows by human researchers. This is followed by the analysis stage, performing a fine-grained interpretation of each file and function with respect to their intended functionality, such as required inputs and outputs, interactions with other modules, and any algorithmic or architectural constraints derived from the source paper. Finally, in the generation stage, the framework synthesizes the entire code base based on the execution order determined earlier, along with the artifacts produced in the previous stages.
@@
Figure 1: (a) PaperCoder overview. The proposed PaperCoder aims to transform given scientific papers (in machine learning domains) into code repositories, which consists of three sequential steps: planning, analyzing, and coding. (b) Code availability. The availability of the code repositories, where blue bars indicate the total number of accepted papers and orange regions represent the subset of papers with officially released code.
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers accepted at top-tier venues in 2024, including NeurIPS, ICML, and ICLR referred to as our proposed Paper2Code benchmark. Furthermore, we also incorporate the recently released PaperBench benchmark in our evaluation suite, enabling the fine-grained evaluations of code implementations with some papers from ICML 2024 [40]. Then, on a battery of tests conducted with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) as well as expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code bases that could meaningfully support human researchers in understanding and reproducing prior work. Specifically, 77% of the generated repositories by PaperCoder are rated as the best, and 85% of human judges report that the generated repositories are indeed helpful. Also, further analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated code bases can be executed, sometimes with only minor modifications (averaging 0.48% of total code lines) in cases where execution errors occur.
@@ 2 Related Work
Large Language Models with Code
Large language models (LLMs) demonstrate impressive capabilities in both text understanding and generation [30, 12, 37]. Recently, their applicability has extended beyond general knowledge tasks into specialized fields such as mathematics, science, and coding, showcasing impressive reasoning and knowledge representation capabilities [34, 43, 41]. Particularly, code-specialized LLMs [10, 9, 17] have garnered significant attention by achieving remarkable performance on various software engineering tasks [45], including software development, software design [36, 15], requirements elicitation [29], and specification formalization [27]. Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of code-specialized LLMs.
@@
LLM-based code generation can be broadly categorized into single-file and multi-file (repository-level) coding. Single-file coding focuses on generating relatively short code snippets to solve isolated tasks, such as programming competition problems or simple coding queries, and was the primary focus in the early stages of LLM-based coding research [21, 6, 1, 14]. As LLMs have advanced in code comprehension, long-context reasoning, and handling complex workflows, research has increasingly shifted toward repository-level coding, where multiple files are generated while jointly considering architectural and functional requirements [23, 18, 25]. Several recent studies have explored this direction [48, 31]. Multi-agent frameworks such as ChatDev leverage LLMs to collaborate through structured dialogues [36], while MetaGPT adopts a waterfall development model with role-specific generation stages [15]. These advancements highlight the growing capability of LLMs to support end-to-end software engineering tasks beyond single-file generation.
LLM-Powered Research
Scientific progress typically involves a cycle of idea generation, hypothesis validation, and experimental execution [33]. LLMs have been applied to various stages of this cycle, including research ideation [2, 20], hypothesis generation [47, 35], and peer review [8, 22, 44], thereby helping researchers overcome existing limitations and accelerate scientific discovery [19, 26, 46]. In computer science and machine learning, where code-based experimentation is fundamental, LLMs have also been utilized to design experiments that enhance existing codebases. Many recent studies focus on refining implementations or optimizing performance, assuming the availability of the original implementation [16, 49, 4, 42]. However, this assumption limits their applicability in cases where source code is unavailable. Reproducibility plays a critical role in scientific discovery, enabling researchers to identify limitations, refine methodologies, and build upon prior work [32]. To evaluate LLMs’ replication capabilities, PaperBench introduced a benchmark in which AI agents attempt to reproduce machine learning papers [40]. While this benchmark emphasizes evaluation, our work addresses a challenge of automatically generating code repositories from scientific papers in the absence of original implementations. To this end, we propose a framework that leverages LLMs for end-to-end research reproducibility. By focusing on repository-level reproduction, our approach expands the scope of LLM-powered automation beyond ideation and hypothesis generation, contributing to a critical yet underexplored aspect of scientific research.
3 Methods
3.1 Problem Definition
Research Repository Generation
In machine learning research, experiments are typically conducted using a code. However, in many cases, researchers do not release their code, making it challenging for others to reproduce and verify the proposed method and experiments [28, 32]. When non-authors attempt to re-implement the code manually, the process is often labor-intensive, time-consuming. To address these challenges, we define the task of Research Repository Generation. The goal of this task is to read a paper and generate a corresponding repository that implements its methods and experiments. We define a repository as a collection of one or more code files necessary for reproducing the methods and experiments described in the paper. By automating this process, research repository generation facilitates the validation and reproducibility of scientific findings. Additionally, new research ideas often build upon prior work. However, when an researcher does not release their code, experimental validation becomes difficult, potentially hindering further advancements. Our approach aims to bridge this gap by enabling researchers to experimentally verify prior work, even when the original implementation is unavailable, thus contributing to a more reproducible and transparent research. In this study, we frame repository generation as a software development problem and assume that a model is responsible for writing the code to create a functional software repository. Given a research paper as input, the model generates a repository containing the code necessary to replicate the paper’s methods and experiments. Formally, we define this as: $M(R)=C$; where $R$ represents the paper, $C$ represents the generated code, and $M$ is a function or model. While $M$ can take various forms, large language models (LLMs) are a representative candidate. A language model functions as a mapping from an input $x$ (text) to an output $y$ (text or code), making it well-suited for this task.
@@ 3.1 Problem Definition
Figure 2: (Left) The naive approach, where a model directly generates code from the paper. (Right) PaperCoder framework, which decomposes the task into three stages: (1) Planning, where a high-level implementation plan is constructed based on the paper’s content, including overall plan, architectural design, logic design, and configuration files; (2) Analyzing, where the plan is translated into detailed file-level specifications; and (3) Coding, where the final codes are generated to implement the paper’s methods and experiments.
3.2 Our Method
We introduce PaperCoder, a novel framework designed specifically for research repository generation that implements the methods and experiments described in a research paper, illustrated in Figure 2. The goal of our work is to model the repository generation process as: $M(R)=C$ where $M$ is the model responsible for transforming the paper into a code repository, $R$ represents the research paper, and $C$ represents the generated code. Inspired by software development methodologies, we adopt a structured approach that mirrors well-established, empirically validated software engineering principles. Specifically, we follow a plan-analysis-implementation workflow to systematically address the problem of research repository generation. To achieve this, we decompose the process into three sequential phases: 1) Planning. 2) Analyzing. 3) Coding. Each phase utilizes task-specialized LLM agents that leveraging a multi-agent scenarios. Formally, we define the repository generation process as: $C=M(P)=M_{\text{code}}(R,P,A)$; where $R$ represents the research paper, $P$ is the plan and $A$ indicates the analysis. Each components are generated as follows: $P=M_{\text{plan}}(R)$, $A=M_{\text{analysis}}(R,P)$ and $C=M_{\text{code}}(R,P,A)$. The following sections provide a detailed explanation of each phase.
@@ 3.2.1 Planning
Providing the research paper as a input to the model and expecting it to generate a complete repository is highly challenging. Research papers are primarily written to document discoveries and persuade readers, rather than to serve as structured input for software development. Consequently, papers often contain supplementary information that, while essential for conveying core concepts, is not directly relevant to implementation. It could be introduces noise, making repository generation difficult and less effective. To address this challenge, instead of solely using the paper as input, we propose decomposing the paper into a structured multi-aspect plan. This approach organizes the key implementation-relevant elements into four distinct components, ensuring that the generated repository is well-structured and aligned with the paper’s methodology. Formally, $M_{\text{plan}}(R)=P=\{o,d,l,g\}$; where $o$ indicates the overall plan, $d$ is the architecture design, $l$ represents the logic design and $g$ is the configuration file. Each output is generated in each phase, and this phase follows a sequential structure, where the output of each stage serves as the input for the next. Note that each phase utilized by the same agent which is the plan agent $M_{\text{plan}}$.
@@
The first step within planning phase, the overall plan, involves summarizing and organizing the core elements necessary to implement the research repository from a high-level perspective. This summary provides an essential conceptual framework that guides subsequent steps clearly. Formally, $M_{\text{plan}}(R)=o$, where $R$ is the research paper and $o$ indicates the overall plan.
@@
The second stage involves architecting the software based on the overall plan generated in the previous stage and the research paper. Designing a well-structured architecture is essential, particularly for software systems where multiple features must interact seamlessly. This stage focuses on identifying the necessary components and defining their relationships to ensure a well-organized and functional repository. To achieve this, we asks to create key artifacts that define the software architecture. The file list is a structured collection of files required for the repository, outlining the software’s modular structure. The class diagram provides a static representation of the system’s data structures and interfaces. Using Unified Modeling Language (UML) notation, a standardized visual language for modeling software systems, we depict classes as rectangles, attributes and methods as lists, and relationships as connecting lines to illustrate how different components interact. The sequence diagram represents the program’s call flow and object interactions dynamically, which uses UML notation to show actors as objects, messages as arrows, and lifelines as dashed lines, visually mapping out how components communicate over time. This structured approach ensures a clear and organized representation of the software’s architecture. By constructing these artifacts, we visually represent the essential components described in the research paper, enabling a more structured and systematic approach to repository generation. This process facilitates better analysis of dependencies and relationships, ensuring that the generated repository aligns with the paper’s core idea. Formally, we define the architecture design process as: $M_{\text{plan}}(R,o)=d$ where $R$ represents the paper, $o$ refers to the previously generated overall plan, and $d$ denotes the resulting architectural design artifacts.
@@
In software development, individual files rarely function in isolation. Instead, they generally exhibit interdependencies through imports and module interactions. For example, if a function A is defined in utils.py and later imported into evaluation.py, utils.py must be implemented before evaluation.py to maintain a correct dependency structure. To account for these dependencies, this stage takes as input the research paper along with the artifacts generated in the previous two stages. It then analyzes the logic of each file and its components, determining the necessary dependencies and the optimal execution order for implementation. As an output, it produces an ordered file list, detailing each file’s role which files should be implemented considering dependencies and its dependencies within the repository. This approach ensures that repository generation considers not only individual file structures but also inter-file communication, facilitating a well-organized and logically coherent implementation. Formally, we define this dependency-aware file ordering process as $M_{\text{plan}}(R,o,d)=l$ where $R$ represents the research paper, $o$ refers to the overall plan, $d$ denotes the architectural specifications, and $l$ is the ordered file list which is the structured output defining the order and role of each file in the repository.
Configuration File Generation
Finally, the configuration file generation step synthesizes all previously established outputs to produce a configuration file (config.yaml) containing hyperparameters and configuration required for model training. This process takes as input the overall plan, architectural design, and ordered file list produced in the previous stages. At this stage, users can review and modify the config.yaml file to identify and correct any missing or incorrectly specified details. For example, users may need to specify the path to a Hugging Face dataset or define the checkpoint storage directory. This step helps reduce hallucinations during the generation process, such as the model producing nonexistent datasets or referencing incorrect file paths. Formally, the configuration generation process is defined as $M_{\text{plan}}(R,o,d,l)=g$ where $R$ represents the research paper, $o$ refers to the overall plan, $d$ denotes the software architecture specifications, $l$ represents the logic design, and $g$ is the resulting configuration file.
3.2.2 Analyzing
While the planning phase primarily focuses on designing the overall repository structure and outlining a high-level roadmap, the analyzing phase delves into the implementation specifics of each individual file. In this phase, the detailed purpose and necessary considerations for each file within the repository are thoroughly analyzed. The outputs generated at this stage clearly specify what each file should achieve and highlight critical factors required for successful implementation. Specifically, the inputs to this analysis phase consist of the original research paper and previously generated artifacts (the overall plan, architecture design, logic designs, and configuration file). The outputs of this phase include file-level analyses documenting precise implementation details, which will later inform the code generation process. Formally, the file-level analysis process is defined as $\{M_{\text{analysis}}(R,P=\{o,d,l,g\},f_{i})=a_{i}\}_{i=1}^{n=|\text{file}|};\text{where}f\in F,a\in A$ where $R$ represents the research paper, $P$ includes four plans, which $o$ refers to the overall plan, $d$ denotes the architectural specifications, $l$ represents the logic design which consists of structured dependencies among files, $g$ is the configuration file, and $f_{i}$ specifies the target file for which the analysis is being conducted. This process is applied to each file individually to generate a $a_{i}$ which is the comprehensive file-level implementation plan.

### v3

@@ Abstract
Despite the rapid growth of machine learning research, corresponding code implementations are often unavailable, making it slow and labor-intensive for researchers to reproduce results and build upon prior work. In the meantime, recent Large Language Models (LLMs) excel at understanding scientific documents and generating high-quality code. Inspired by this, we introduce PaperCoder, a multi-agent LLM framework that transforms machine learning papers into functional code repositories. PaperCoder operates in three stages: planning, where it constructs a high-level roadmap, designs the system architecture with diagrams, identifies file dependencies, and generates configuration files; analysis, which focuses on interpreting implementation-specific details; and generation, where modular, dependency-aware code is produced. Moreover, each phase is instantiated through a set of specialized agents designed to collaborate effectively across the pipeline. We then evaluate PaperCoder on generating code implementations from machine learning papers based on both model-based and human evaluations, particularly from the authors of those papers, with author-released repositories as ground truth if available. Our results demonstrate the effectiveness of PaperCoder in creating high-quality, faithful implementations. Furthermore, it consistently shows strengths in the recently released PaperBench benchmark, surpassing strong baselines by substantial margins. Code is available at: https://github.com/going-doer/Paper2Code.
@@ 1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge [6, 3, 30]. However, across many disciplines, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only average 19.5% of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of scientific innovation.
@@
To answer this question, we introduce PaperCoder, a multi-agent LLM-powered framework, designed to automatically generate faithful code repositories in machine learning directly from and contextualized with research papers, which differs from prior work that requires partial implementations from human inputs. Specifically, PaperCoder aims to emulate the typical life cycle of human developers and researchers in writing the repository-level code, by decomposing the task into three structured stages: planning, analysis, and generation. First, during the planning stage, the proposed framework constructs a high-level roadmap to identify core components to implement, draws the overall system architecture with class and sequence diagrams to model structural relationships between modules, identifies file dependencies with their execution orders to guide correct build and execution flows, and generates configuration files to enable flexible customization of experimental workflows by human researchers. This is followed by the analysis stage, performing a fine-grained interpretation of each file and function with respect to their intended functionality, such as required inputs and outputs, interactions with other modules, and any algorithmic or architectural constraints derived from the source paper. Finally, in the generation stage, the framework synthesizes the entire code base based on the execution order determined earlier, along with the artifacts produced in the previous stages.
@@
Figure 1: (a) PaperCoder overview. The proposed PaperCoder aims to transform given scientific papers (in machine learning domains) into code repositories, which consists of three sequential steps: planning, analysis, and coding. (b) Code availability, where blue bars indicate the total number of accepted papers and orange regions show those with officially released code (See Appendix B.1 for details on the availability calculation).
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers accepted at top-tier venues in 2024, including ICLR, ICML, and NeurIPS referred to as our proposed Paper2Code benchmark (in short, Paper2CodeBench). Also, we incorporate the recent PaperBench benchmark [38] in our evaluation suite, enabling fine-grained evaluations of code implementations with some papers from ICML 2024. Then, on a battery of tests conducted not only with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) but also with expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code repositories that could meaningfully support human researchers in understanding and reproducing prior work. Specifically, 88% of the generated repositories by PaperCoder are rated as the best over baselines, and 92% of human judges report that the generated repositories are indeed helpful. Also, further analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated codebases can be executed, sometimes with only minor modifications (averaging 0.81% of total code lines) in cases where execution errors occur.
@@ 2 Related Work
Large Language Models for Code
LLMs have shown impressive capabilities in text understanding and generation [28, 10, 35], which have been widely utilized within specialized domains (beyond general tasks), such as mathematics, science, and coding [32, 43, 40]. Particularly, code-specialized LLMs [9, 8, 15] have received significant attention thanks to remarkable performance on various software engineering tasks [45], including software design and development [34, 13], requirements elicitation [27], and formal specification generation [25]. Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
@@
Early work on LLM-based code generation typically focuses on single-file tasks, whose objective is to generate relatively short code snippets to solve isolated tasks, such as (algorithmic-level) programming competition problems or toy-level programming queries [19, 5, 1, 12]. However, as LLMs have advanced in comprehending and generating code with the ability to perform long-context reasoning, recent studies have increasingly shifted their attention toward more challenging repository-level coding tasks, which involve generating multi-file repositories that jointly account for architectural design, modular structure, and inter-file dependencies [21, 16, 39]. In particular, several recent efforts explore this emerging paradigm [50, 29], often adopting multi-agent or role-based frameworks to emulate realistic development workflows. For instance, ChatDev instantiates LLMs into role-playing agents that collaborate through structured dialogues [34], while MetaGPT implements a waterfall-style development pipeline with specialized agents for handling different stages of software engineering [13]. Beyond prior work, we explore the underexplored yet challenging task of transforming full-length, complex papers into repository-level implementations.
LLM-Powered Scientific Research
LLMs have been increasingly adopted to support every stage of the scientific process from ideation to experimental validation [31, 2, 18, 48, 33, 7, 20, 44]; thereby, helping researchers overcome existing challenges and ultimately accelerate scientific discovery [17, 24, 47]. Specifically, in fields such as computer science and machine learning (where code-based experimentation is central), LLMs have been used to design, refine, and extend code implementations. Yet, many recent efforts in this space assume access to and build on top of the original codebase [14, 46, 4, 41], which significantly limits their applicability in real-world scenarios since such implementations are oftentimes unavailable (Figure 1). To address this, concurrent to our work, Starace et al. [38] introduces a benchmark dataset called PaperBench, evaluating the capability of existing agentic AI systems in reproducing papers with fine-grained metrics. Notably, on top of PaperBench (which emphasizes evaluation), we further complement and extend this line by focusing on methodological aspects of how to transform scientific papers into repository-level code implementations.
3 Method
In this section, we start with describing the task of repository-level code generation from machine learning papers, and propose PaperCoder, a multi-agent, multi-stage framework designed to tackle it.
3.1 Repository-Level Code Generation from Scientific Papers in Machine Learning
The goal of our repository-level code generation task is to automatically produce a code repository that faithfully implements the methods and experiments described in machine learning papers (especially for cases where authors do not release their code), to support reproducibility and accelerate scientific progress [26, 30]. Formally, we define this task as a function (or a model) $M$ that maps a research paper $R$ to a corresponding code repository $C$, as follows: $M(R)=C$. Here, $C$ is composed of multiple interdependent files $\{c_{1},c_{2},...,c_{n}\}$, each responsible for implementing different components of the methods and experiments described in $R$, but together they should form a cohesive pipeline.
It is worth noting that the most straightforward approach to instantiating $M$ is to instruct the LLM to generate the entire code repository, conditioned on the given paper, as follows: $M(R):=\texttt{LLM}(\mathcal{T}(R))$, where $\mathcal{T}$ is the prompt template that specifies the intended behavior of the LLM for the target task (including task descriptions, detailed instructions, and any other relevant context). Yet, generating a complete, modular, and faithful repository (spanning multiple files) in a single pass is extremely challenging, even for powerful LLMs, due to the inherent complexity of scientific papers and their corresponding implementations, the long-context limitations of current models, and the difficulty in maintaining consistent global structure and cross-file dependencies. Therefore, we propose a multi-stage and multi-agent framework that decomposes the overall task into smaller, manageable subtasks, each handled by a specialized agent tailored to a specific aspect of paper-to-code transformation.
@@ 3.1 Repository-Level Code Generation from Scientific Papers in Machine Learning
Figure 2: (Left) The naive approach, which directly generates an entire code repository from a paper. (Right) Our PaperCoder framework, which is operationalized by decomposing the task into three stages: (1) Planning, where a high-level implementation plan is constructed from the paper, including overall plan, architectural design, logic design, and configuration file; (2) Analysis, where the plan is translated into detailed file-level specifications; and (3) Coding, where the final codes are generated to implement the methods and experiments of the paper.
3.2 PaperCoder: LLM-Powered Multi-Agent Framework for Paper-to-Code Transformation
We now introduce PaperCoder, a structured, multi-agent framework for generating code repositories directly from machine learning papers (without access to pre-existing artifacts or implementations, such as skeleton code). Specifically, inspired by typical software development workflows, PaperCoder decomposes the task into three coordinated stages: Planning, Analysis, and Coding, each orchestrated by specialized LLM agents. Formally, given a paper $R$, the overall process can be defined as follows:
$\text{{Planning: }}P=M_{\text{plan}}(R),\quad\text{{Analysis: }}A=M_{\text{analysis}}(R,P),\quad\text{{Coding: }}C=M_{\text{code}}(R,P,A),$
where $P$, AA, and $C$ represent the high-level implementation plan, the detailed function-level analysis, and the final code repository, respectively. The overall pipeline of PaperCoder is shown in Figure 2.
@@ 3.2.1 Planning
It is worth noting that, in contrast to implementation specifications designed explicitly for software development, scientific papers are written to communicate ideas and findings to humans. As a result, they often contain high-level motivations, persuasive narratives, and auxiliary details that are crucial for human understanding but noisy, loosely specified, or ambiguous from a software engineering perspective. To mitigate this challenge, we introduce a structured planning phase that transforms unstructured textual content into implementation-level abstractions. Also, we decompose the planning process into four sequential subcomponents (to simplify the task and reduce cognitive and computational load of LLM-powered agents at each step): 1) overall plan, 2) architecture design, 3) logic design, and 4) configuration generation. Formally, we define this as: $M_{\text{plan}}(R)\rightarrow P=\{o,d,l,g\}$, where $o$ is the overall plan, $d$ is the architecture design, $l$ is the logic design, and $g$ is the configuration file, with each stage using the outputs of the previous ones as contextual input. We then describe how each subcomponent is instantiated through specific objectives and instructions below.
@@
The first step is to extract a high-level summary of the core components and functionalities described throughout the paper, to identify the specific methods and experiments to be implemented. In other words, this high-level overview includes model components, training objectives, data processing steps, and evaluation protocols (distributed across the entire paper), which can form the foundation for all subsequent steps, formalized as follows: $M_{\text{plan}}^{(1)}(R):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(1)}(R))\rightarrow o$.
@@
Based on the extracted overall plan alongside the input paper, the next step is to define the repository-level architecture, which includes identifying files, organizing them into modules, and defining their relationships, to ensure a coherent and maintainable structure. Specifically, the LLM-powered agent is prompted to generate a file list, which outlines the overall file structure of the repository; a class diagram, which details static representations of files (such as core classes and their attributes); and a sequence diagram, which models the dynamic interactions. Formally, similar to overall plan, this process can be defined as follows: $M_{\text{plan}}^{(2)}(R,o):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(2)}(R,o))\rightarrow d$.
@@
While the previous architecture design focuses on what to build, the logic design phase specifies how these components should be instantiated in practice by considering their dependencies in terms of overall execution flow. In other words, its primary goal is to translate abstract component-level relationships from the architecture design into an actionable execution order for the entire implementation. This step is crucial because individual modules often depend on shared utilities, configurations, or data loaders that are defined in other parts of the repository, and without an explicitly defined execution order, the code generation can result in failure or inconsistency (e.g., generating file B before file A when B imports modules from A). To address this, the logic design stage not only produces an ordered file list that dictates the sequence in which the files should be implemented and executed, but also further elaborates on the logic within each file; thereby, providing more fine-grained specifications. Formally, $M_{\text{plan}}^{(3)}(R,o,d):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(3)}(R,o,d))\rightarrow l$.
Configuration Generation
In the last stage of planning, PaperCoder synthesizes a configuration file (namely, config.yaml) that includes key hyperparameters, model settings, and other runtime options based on prior outputs alongside the given paper. We note that, in addition to grounding the code generation process with the explicit configuration details, it enables researchers to easily review and adjust experimental configurations without modifying the source code. Formally, $M_{\text{plan}}^{(4)}(R,o,d,l):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(4)}(R,o,d,l))\rightarrow g$. We provide prompts used to elicit each planning output in Appendix E.
3.2.2 Analysis
Following the planning stage, which defines the overall structure and execution flow of the repository, the analysis phase focuses on interpreting and specifying the implementation-level details for modules within each file. In other words, unlike planning that answers what components to build and how they relate, this phase addresses the question of how each component should be operationalized and concretely implemented at the file level, which includes the definition of functional goals, input-output behaviors, intra- and inter-file dependencies, and algorithmic specifications derived from the original paper. Specifically, given the input paper $R$ and planning outputs $P=\{o,d,l,g\}$, the analysis agent iteratively processes each file $f_{i}$ (identified during planning) and generates a detailed analysis $a_{i}$ describing what needs to be implemented in that file. Formally, $\{M_{\text{analysis}}(R,P,f_{i})\}_{i=1}^{n=|F|}$ where $M_{\text{analysis}}(R,P,f_{i}):=\texttt{LLM}(\mathcal{T}_{\text{analysis}}(R,P,f_{i}))\rightarrow a_{i}$, with $F$ as the set of identified files, e.g., $f_{i}\in F$.